OpenStudy (fouzberzerk):

What is the value of b? b/-3 = 1.5 It's a fraction.

OpenStudy (fouzberzerk):

@kimbllse777

OpenStudy (tkhunny):

You have shown no work of your own. Please give it a go.

OpenStudy (fouzberzerk):

Verifying my own answers is not cheating.

OpenStudy (tkhunny):

I said nothing of verifying anything. I said only that you had SHOWN no work of your own. Maybe you did it already. Maybe you didn't. This is not the question. SHOW YOUR WORK.

OpenStudy (fouzberzerk):

@tkhunny Yeah, showing my work in an online assessment is so important. Thanks for enlitening me oh great one.

OpenStudy (anonymous):

-3×1.5=-4.5

OpenStudy (fouzberzerk):

Alright, I have another question.

OpenStudy (anonymous):

K

OpenStudy (fouzberzerk):

olve the equation for x. 7x = –2y A. x = - 2y/7 B. x = -14y C. x = - 7/2y D. x = - 9y

OpenStudy (fouzberzerk):

A and C are both fractions.

OpenStudy (fouzberzerk):

I think it's B.

OpenStudy (anonymous):

B is correct

OpenStudy (fouzberzerk):

Thanks.

OpenStudy (fouzberzerk):

"Issued a warning" top keks/10

OpenStudy (fouzberzerk):

Which sets of numbers are closed under multiplication? Choose all answers that are correct. A. whole numbers B. natural numbers C. negative integers D. integers

OpenStudy (fouzberzerk):

Would you bugger off @tkhunny?

OpenStudy (fouzberzerk):

@kimbllse777

OpenStudy (anonymous):

I'm not really sure. I haven't really learned that much about closure

OpenStudy (tkhunny):

What do you think,@fouzberzerk What is closure and how do you demonstrate it?